Jessica has two bags of nuts; one with 76
% peanuts and the other with 12% peanuts. To make and 11lb bag of 47% peanuts, how much of each should Jessica use.

Answer by fractalier(6550)

Call the amount of 76% peanuts, x.
The amount of 12% would then be 11-x.
The setup looks like this
76(x)+12(11-x)= 11(47)
Now solve this.
76x + 132 - 12x = 517
64x = 385
x = 6.016 lbs of 76%
11-x = 4.984 lbs of 12%